OpenStudy (anonymous):

If f(x) =1/2 x + 2 and g(x) = −3x − 1, find g[f(4)].

OpenStudy (anonymous):

Hey again. for this problem find f(4) by pluging in 4 for x in f(x). Then everywhere you find x in g(x), plug in f(4).

OpenStudy (anonymous):

dont get it

OpenStudy (anonymous):

do you understand how to find f(4)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay. f(x) =1/2 x + 2. so f(4) is replace x with the value 4. f(4) = 1/2 (4) +2 = 4

OpenStudy (anonymous):

is tht the answer

OpenStudy (anonymous):

no once you have that. replace x in g(x) with 4

OpenStudy (anonymous):

g(4) = -3(4) -1

OpenStudy (anonymous):

brandon 77 ure the pellet

OpenStudy (anonymous):

????

OpenStudy (anonymous):

\[gof(x)=-3(\frac{ 1 }{ 2 }x+2)-1\] \[gof(x)=-\frac{ 3x }{ 2 }-7\] Put 4 where 4 is and solve.

OpenStudy (anonymous):

Put 4 where x is, I mean.

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

-13